OpenStudy (anonymous):

find the second derivative of x^n when n is greater than or equal to 2

OpenStudy (anonymous):

power rule

OpenStudy (anonymous):

but what is throwing me off is the greater than sign

OpenStudy (anonymous):

just redundant information

OpenStudy (anonymous):

if x=1,then second or higher derivatives=0

geerky42 (geerky42):

You mean n=1?

OpenStudy (campbell_st):

start with this if you were given \[y = x^n\] what would you write for the 1st derivative...?

OpenStudy (anonymous):

nx^(n-1)

OpenStudy (campbell_st):

great and n is just a constant so now what is the 2nd derivative...?

OpenStudy (campbell_st):

of just differentiate you last answer

OpenStudy (campbell_st):

so just apply the same method to find the 2nd derivative... \[y'= x^{n - 1}\] then multiply the answer by n

OpenStudy (anonymous):

wouldn't it just be (n-1)(n-2)x^(n-2)

geerky42 (geerky42):

\[x^n\\~\\~\\\dfrac{d}{dx}x^n = nx^{n-1}\]Let \(m = n-1\) So we have \(\dfrac{d}{dx}x^n = nx^{n-1}=nx^{m}\) Now taking derivative again: \(\dfrac{d}{dx}n~x^m = n~m~x^{m-1} = n(n-1)x^{(n-1)-1} = \boxed{n(n-1)x^{n-2}}\)